# Evaluating diversion and treatment policies for opioid use disorder

### v3

@@ 3 Results and Discussion
In terms of societal costs, Table 4 shows that statistically significant global total expected societal cost savings through the Year 2032, compared to the base model, is between \$68.35 million at a (0, 30, 0) OD-focused policy to \$1.06 billion at the policy-mix of (60, 80, 60). Additional fixed and variable cumulative policy implementation costs could be subtracted from the expected societal cost savings to get a total savings estimate. At and above the (20, 22, 0) AD-focused policy, the (0, 30, 0) OD-only policy, the (0, 22, 20) CM-focused policy, or a policy-mix of (20, 40, 20), we see that the cost savings from decreased opioid use, hospital encounters, and opioid-related arrests outweigh the added costs of additional OUD treatment. Decreased opioid use and opioid-related arrests increase the quality of life and productivity of the community (Florence et al.,, 2021). Therefore, jurisdictions that invest in additional treatment capacity and divert individuals to treatment can expect to improve residents’ lives, increase the active workforce, and reduce incarceration.
Figure 3 shows the upper 95% PI usage of the CJS due to opioid-related arrests, OUD treatment usage, and hospital usage due to an opioid encounter for each year. The upper 95% PI can be interpreted as the predicted maximum resource usage at a given time. If the capacity of a resource is below the predicted maximum usage, there is a risk of resource shortages at some point in the year. Figure 3(a) shows that implementing an AD-focused policy reduces the number of individuals arrested for an opioid-related offense in the CJS. For example, implementing policy-mixes with AD at 60%, i.e., (60, 22, 0) and (60, 80, 60), reduces the CJS maximum population by a magnitude of 35. Figure 3(b) shows that for any additional policy, usage of OUD treatment goes up and compounds with more/higher policies. For example, implementing the policy-mix of (60, 80, 60) requires 1.3 to 1.8 times the amount of OUD treatment capacity than the base model. Additionally, we observe a positive trend in the demand for OUD treatment for all policies, including the base model. Therefore, even if jurisdictions implement no treatment policies, due to the nature of the opioid crisis, demand for OUD treatment capacity is predicted to increase on average by 2% to 5% each year. Increasing OUD treatment capacity is important for jurisdictions to consider since most treatment centers are currently at or near capacity (Jones et al.,, 2015). Additionally, expansion of OUD treatment is needed when implementing a treatment policy that redirects individuals to facilities already at or near capacity. Lastly, Figure 3(c) shows that hospital usage does not significantly differ in any treatment policy.
@@
Table 5 shows a statistically significant decrease in re-arrests for any AD and CM policy. For example, Policy (60, 22, 0) reduces the re-arrest rate by 41% compared to the base model in 2032. We also see statistically significant decreases in re-hospitalization when implementing the OD policy at any level. All policies increase the number of individuals in and out of OUD treatment. Therefore, OUD treatment policies may help reduce “frequent fliers” in the CJS and hospital systems while redirecting them to treatment resources. Additional results are in Appendix C.
Table 5: Simulated mean $\pm$ standard error (p-value) of Year 2032 opioid-related re-arrest, hospitalization, and treatment re-start rates for all scenarios
@@
Base Model 0, 22, 0 5.19 $\pm$ 0.022 5.17 $\pm$ 0.02 5.24 $\pm$ 0.019 0.77 $\pm$ 0.014 0.78 $\pm$ 0.014 0.78 $\pm$ 0.014 0.45 $\pm$ 0.006 0.45 $\pm$ 0.005 0.45 $\pm$ 0.005
Arrest Diversion 20, 22, 0 5.22 $\pm$ 0.021 5.19 $\pm$ 0.02 5.23 $\pm$ 0.019 0.66 $\pm$ 0.013∗∗ 0.65 $\pm$ 0.012∗∗ 0.66 $\pm$ 0.012∗∗ 0.5 $\pm$ 0.006∗∗ 0.51 $\pm$ 0.006∗∗ 0.51 $\pm$ 0.005∗∗
40, 22, 0 5.22 $\pm$ 0.021 5.18 $\pm$ 0.021 5.21 $\pm$ 0.02 0.57 $\pm$ 0.013∗∗ 0.54 $\pm$ 0.012∗∗ 0.57 $\pm$ 0.012∗∗ 0.56 $\pm$ 0.006∗∗ 0.55 $\pm$ 0.006∗∗ 0.55 $\pm$ 0.005∗∗
60, 22, 0 5.17 $\pm$ 0.02 5.16 $\pm$ 0.021 5.19 $\pm$ 0.019 0.44 $\pm$ 0.011∗∗ 0.44 $\pm$ 0.011∗∗ 0.46 $\pm$ 0.011∗∗ 0.61 $\pm$ 0.006∗∗ 0.59 $\pm$ 0.006∗∗ 0.6 $\pm$ 0.006∗∗
80, 22, 0 5.18 $\pm$ 0.022 5.13 $\pm$ 0.02 5.16 $\pm$ 0.019 0.34 $\pm$ 0.009∗∗ 0.34 $\pm$ 0.009∗∗ 0.34 $\pm$ 0.009∗∗ 0.64 $\pm$ 0.006∗∗ 0.64 $\pm$ 0.006∗∗ 0.64 $\pm$ 0.006∗∗
100, 22, 0 5.17 $\pm$ 0.021 5.13 $\pm$ 0.02 5.12 $\pm$ 0.019∗∗ 0.24 $\pm$ 0.008∗∗ 0.22 $\pm$ 0.007∗∗ 0.23 $\pm$ 0.007∗∗ 0.7 $\pm$ 0.007∗∗ 0.7 $\pm$ 0.006∗∗ 0.69 $\pm$ 0.006∗∗
Overdose Diversion 0, 30, 0 4.81 $\pm$ 0.019∗∗ 4.79 $\pm$ 0.019∗∗ 4.84 $\pm$ 0.019∗∗ 0.75 $\pm$ 0.015 0.77 $\pm$ 0.014 0.78 $\pm$ 0.013 0.58 $\pm$ 0.006∗∗ 0.58 $\pm$ 0.006∗∗ 0.57 $\pm$ 0.006∗∗
0, 45, 0 4.06 $\pm$ 0.019∗∗ 4.03 $\pm$ 0.019∗∗ 4.06 $\pm$ 0.017∗∗ 0.78 $\pm$ 0.015 0.78 $\pm$ 0.015 0.77 $\pm$ 0.013 0.8 $\pm$ 0.007∗∗ 0.8 $\pm$ 0.007∗∗ 0.81 $\pm$ 0.006∗∗
0, 60, 0 3.34 $\pm$ 0.016∗∗ 3.31 $\pm$ 0.016∗∗ 3.31 $\pm$ 0.015∗∗ 0.77 $\pm$ 0.014 0.75 $\pm$ 0.014 0.77 $\pm$ 0.014 1.05 $\pm$ 0.007∗∗ 1.04 $\pm$ 0.007∗∗ 1.04 $\pm$ 0.007∗∗
0, 75, 0 2.6 $\pm$ 0.015∗∗ 2.61 $\pm$ 0.014∗∗ 2.6 $\pm$ 0.013∗∗ 0.75 $\pm$ 0.015 0.76 $\pm$ 0.014 0.74 $\pm$ 0.014 1.28 $\pm$ 0.008∗∗ 1.29 $\pm$ 0.008∗∗ 1.28 $\pm$ 0.007∗∗
0, 90, 0 1.88 $\pm$ 0.012∗∗ 1.87 $\pm$ 0.012∗∗ 1.86 $\pm$ 0.011∗∗ 0.75 $\pm$ 0.014 0.76 $\pm$ 0.014 0.74 $\pm$ 0.014 1.5 $\pm$ 0.008∗∗ 1.5 $\pm$ 0.008∗∗ 1.49 $\pm$ 0.007∗∗
Case Management 0, 22, 20 5.21 $\pm$ 0.021 5.19 $\pm$ 0.021 5.21 $\pm$ 0.019 0.67 $\pm$ 0.014∗∗ 0.65 $\pm$ 0.013∗∗ 0.68 $\pm$ 0.013∗∗ 0.53 $\pm$ 0.006∗∗ 0.54 $\pm$ 0.006∗∗ 0.54 $\pm$ 0.005∗∗
0, 22, 40 5.21 $\pm$ 0.022 5.2 $\pm$ 0.019 5.2 $\pm$ 0.02 0.53 $\pm$ 0.012∗∗ 0.54 $\pm$ 0.011∗∗ 0.55 $\pm$ 0.012∗∗ 0.67 $\pm$ 0.007∗∗ 0.66 $\pm$ 0.006∗∗ 0.66 $\pm$ 0.006∗∗
0, 22, 60 5.19 $\pm$ 0.022 5.14 $\pm$ 0.02 5.21 $\pm$ 0.019 0.43 $\pm$ 0.011∗∗ 0.41 $\pm$ 0.01∗∗ 0.42 $\pm$ 0.01∗∗ 0.84 $\pm$ 0.007∗∗ 0.84 $\pm$ 0.007∗∗ 0.84 $\pm$ 0.006∗∗
0, 22, 80 5.15 $\pm$ 0.021 5.18 $\pm$ 0.021 5.16 $\pm$ 0.019 0.3 $\pm$ 0.01∗∗ 0.31 $\pm$ 0.009∗∗ 0.32 $\pm$ 0.008∗∗ 1.03 $\pm$ 0.007∗∗ 1.02 $\pm$ 0.006∗∗ 1.03 $\pm$ 0.007∗∗
0, 22, 100 5.16 $\pm$ 0.021 5.12 $\pm$ 0.021 5.14 $\pm$ 0.019∗∗ 0.2 $\pm$ 0.007∗∗ 0.2 $\pm$ 0.007∗∗ 0.19 $\pm$ 0.007∗∗ 1.26 $\pm$ 0.008∗∗ 1.23 $\pm$ 0.007∗∗ 1.23 $\pm$ 0.007∗∗
Policy Mix 20, 40, 20 4.31 $\pm$ 0.02∗∗ 4.25 $\pm$ 0.018∗∗ 4.31 $\pm$ 0.018∗∗ 0.58 $\pm$ 0.013∗∗ 0.55 $\pm$ 0.012∗∗ 0.58 $\pm$ 0.012∗∗ 0.86 $\pm$ 0.007∗∗ 0.86 $\pm$ 0.007∗∗ 0.86 $\pm$ 0.006∗∗
40, 60, 40 3.3 $\pm$ 0.017∗∗ 3.3 $\pm$ 0.015∗∗ 3.32 $\pm$ 0.016∗∗ 0.42 $\pm$ 0.011∗∗ 0.43 $\pm$ 0.011∗∗ 0.41 $\pm$ 0.01∗∗ 1.32 $\pm$ 0.008∗∗ 1.32 $\pm$ 0.008∗∗ 1.31 $\pm$ 0.007∗∗
60, 80, 60 2.31 $\pm$ 0.014∗∗ 2.28 $\pm$ 0.013∗∗ 2.33 $\pm$ 0.013∗∗ 0.31 $\pm$ 0.01∗∗ 0.29 $\pm$ 0.008∗∗ 0.29 $\pm$ 0.008∗∗ 1.76 $\pm$ 0.008∗∗ 1.75 $\pm$ 0.008∗∗ 1.75 $\pm$ 0.008∗∗
@@ 3.1 Policy Effect Comparison
To compare policies and assess marginal returns, we evaluate the comparative effects between our three policies AD, OD, and CM, and each of the simulation main outputs: overdose-related deaths, number of hospital encounters, number of non-diverted arrests, number of OUD treatment, and number of individuals in the active state at the end of a given year. We use Ordinary Least Squares (OLS) regression, where the policies are variables against each output in years 2023, 2027, and 2032. Table 6 shows the regression coefficients against each policy and the model outputs across 600 replications for each of the scenarios listed in Tables 4 and 5.
As shown in Table 6, all of the policies are consistent in their direction of influence over model outputs. This means the policies are all negative or all positive for a given year and model output. This means that the policies are complementary in nature and that increasing any policy will likely improve model outputs. We also see that the OD policy has the highest magnitude of coefficients, followed by the CM policy. Therefore, assuming any added implementation costs for increasing any policy by 1% are equal and one could only improve a single policy, an improvement in OD would likely have the largest impact, followed by the CM policy and then the AD policy. This makes sense, given the model input parameters, since a larger number of individuals experience a hospital encounter than with the CJS It also makes sense that the CM policy has a larger impact than the AD policy since it can engage any individual in the CJS with past opioid misuse rather than just individuals who are in the CJS for an opioid-related crime.
3.2 Model Sensitivity
To assess model sensitivity, we used the (60,80,60) scenario and sampled 1024 Sobol sets across 41 parameters, ranging each parameter $\pm$ 5% of their base model value. We ran each set with three replications for a total of 3,072 simulation runs. The sensitivity analysis in Appendix D suggests that this paper’s conclusions are not sensitive to model inputs, since the sensitivity analysis shows that the longer the model is run, the less sensitive the model is to the starting population. Additionally, the relationship between the model inputs and outputs is expected due to the model structure. The effect size of the outputs decreases or does not change as policies are implemented. The exception is Arc (3), where its effect size on opioid-related death increases after the OD is implemented. Overall, care should be taken when estimating distributions for arcs (2)-(6) and (G) since these have the largest effect on base model outputs, particularly $\mu$ for these distributions. Overall, the sensitivity analysis suggests that changes to the model parameters would not affect the conclusions of this paper between the base model and treatment policy scenarios regarding opioid-related arrests, hospital encounters, treatments, and active use.
@@ 3.2 Model Sensitivity
2023 Opioid-related death 1244.2 ($<$ 0.001) -0.22 ($<$ 0.001) -0.04 ($<$ 0.001) -0.02 ($<$ 0.001)
2023 Opioid-related arrest 8413.14 ($<$ 0.001) -1.42 ($<$ 0.001) -0.3 ($<$ 0.001) -0.08 ($<$ 0.001)
2023 Hospital encounters 30853.45 ($<$ 0.001) -5.2 ($<$ 0.001) -1.08 ($<$ 0.001) -0.49 ($<$ 0.001)
2023 Treatment 26377.29 ($<$ 0.001) 131.24 ($<$ 0.001) 37.74 ($<$ 0.001) 12.53 ($<$ 0.001)
2023 Active use 232902.93 ($<$ 0.001) -39.59 ($<$ 0.001) -7.84 ($<$ 0.001) -3.13 ($<$ 0.001)
2027 Opioid-related death 1640.66 ($<$ 0.001) -0.31 ($<$ 0.001) -0.05 ($<$ 0.001) -0.09 ($<$ 0.001)
2027 Opioid-related arrest 10983.49 ($<$ 0.001) -2.04 ($<$ 0.001) -0.39 ($<$ 0.001) -0.56 ($<$ 0.001)
2027 Hospital encounters 39897.67 ($<$ 0.001) -7.2 ($<$ 0.001) -1.42 ($<$ 0.001) -2.27 ($<$ 0.001)
2027 Treatment 34157.2 ($<$ 0.001) 210.94 ($<$ 0.001) 59.15 ($<$ 0.001) 64.66 ($<$ 0.001)
2027 Active use 306554.25 ($<$ 0.001) -57.3 ($<$ 0.001) -10.84 ($<$ 0.001) -15.79 ($<$ 0.001)
2032 Opioid-related death 2177.5 ($<$ 0.001) -0.43 ($<$ 0.001) -0.07 ($<$ 0.001) -0.16 ($<$ 0.001)
2032 Opioid-related arrest 14442.76 ($<$ 0.001) -2.76 ($<$ 0.001) -0.52 ($<$ 0.001) -1.01 ($<$ 0.001)
2032 Hospital encounters 52058.64 ($<$ 0.001) -9.61 ($<$ 0.001) -1.86 ($<$ 0.001) -3.78 ($<$ 0.001)
2032 Treatment 44668.87 ($<$ 0.001) 318.91 ($<$ 0.001) 88.18 ($<$ 0.001) 134.31 ($<$ 0.001)
2032 Active use 405913.14 ($<$ 0.001) -78.43 ($<$ 0.001) -14.51 ($<$ 0.001) -28.22 ($<$ 0.001)

### v4

@@ 3.2 Societal Costs
In terms of societal costs, Table 4 shows statistically significant 10-year societal cost savings through the Year 2032, compared to the base model, is between \$38.65 million at a (0, 22, 20) CM-focused policy to \$584 million at the policy-mix of (60, 80, 60). At and above the (20, 22, 0) AD-focused policy, the (0, 45, 0) OD-only policy, the (0, 22, 20) CM-focused policy, or a policy-mix of (20, 40, 20), we see that the cost savings from decreased opioid use, hospital encounters, and opioid-related arrests outweigh the added costs of additional OUD treatment. Decreased opioid use and opioid-related arrests increase the quality of life and productivity of the community (Florence et al.,, 2021). Therefore, jurisdictions that invest in additional treatment capacity and divert individuals to treatment can expect to improve residents’ lives, increase the active workforce, and reduce incarceration.
Since the reported societal cost savings omits implementation costs, one could reduce the reported savings in this study by incorporating a fixed amount to initiate the program and yearly operating costs. Operating costs might include the costs of OUD treatment, program coordinators, additional OUD treatment staff, and overhead for OUD facilities. To provide some context, we estimate that implementing an intervention policy costs approximately 10 times less than the model’s estimated cost reduction. For example, in 2016, the city of Madison, WI received \$700,000 in Federal funding to implement and study an arrest diversion program for two years (City of Madison,, 2016). MPD received \$$\$1.2$ million of funding in 2019 for two additional years to expand the arrest diversion program to cover the entirety of Dane County and incorporate additional pathways to treatment (City of Madison,, 2019). Extrapolating the program implementation cost at $1.2$ million for two years, we estimate the implementation would cost \$6 million over 10 years. By comparison, our case study of Dane County, where Madison is located, estimates that successfully diverting 20% of eligible individuals to arrest diversion would yield a mean cost savings of \$$\$102.61$ million over ten years or $10.26$ times the cost of Madison AD policy implementation. We assume that other interventions would be comparable in the ratio of implementation costs to societal costs.
3.3 CJS, OUD Treatment, and Hospital Usage
Figure 3 shows the upper 95% PI usage of the CJS due to opioid-related arrests, OUD treatment usage, and hospital usage due to an opioid encounter for each year.
@@ 3.3 CJS, OUD Treatment, and Hospital Usage
The upper 95% PI can be interpreted as the predicted maximum resource usage at a given time. If the capacity of a resource is below the predicted maximum usage, there is a risk of resource shortages at some point in the year. Figure 3(a) shows that implementing an AD-focused policy reduces the number of individuals arrested for an opioid-related offense in the CJS. For example, implementing policy-mixes with AD at 60%, i.e., (60, 22, 0) and (60, 80, 60), reduces the CJS maximum population by a magnitude of 35. Figure 3(b) shows that for any additional policy, usage of OUD treatment goes up and compounds with more/higher policies. For example, implementing the policy-mix of (60, 80, 60) requires 1.3 to 1.8 times the amount of OUD treatment capacity than the base model. Additionally, we observe a positive trend in the demand for OUD treatment for all policies, including the base model. Therefore, even if jurisdictions implement no treatment policies, due to the nature of the opioid crisis, demand for OUD treatment capacity is predicted to increase on average by 2% to 5% each year. Increasing OUD treatment capacity is important for jurisdictions to consider since most treatment centers are currently at or near capacity (Jones et al.,, 2015). Additionally, expansion of OUD treatment is needed when implementing a treatment policy that redirects individuals to facilities already at or near capacity. Lastly, Figure 3(c) shows that hospital usage does not significantly differ in any treatment policy.
3.4 Re-Hospitalization, Re-Arrest, and OUD Treatment Re-Start Rates
@@
Table 5 shows a statistically significant decrease in re-arrests for any AD and CM policy. For example, the AD-focused policy (60, 22, 0) reduces the re-arrest rate by 0.32 compared to the base model in 2032. We also see statistically significant decreases in re-hospitalization when implementing the OD policy at any level. All policies increase the number of individuals in and out of OUD treatment. Therefore, OUD treatment policies may help reduce “frequent fliers” in the CJS and hospital systems while redirecting them to treatment resources. Additional results are in Appendix C.
Table 5: Simulated mean $\pm$ standard error (p-value) of Years’ 2023, 2027, and 2032 opioid-related re-arrest, hospitalization, and treatment re-start rates for all scenarios
@@
Base Model 0, 22, 0 5.18 $\pm$ 0.02 5.19 $\pm$ 0.02 5.22 $\pm$ 0.019 0.76 $\pm$ 0.014 0.8 $\pm$ 0.014 0.76 $\pm$ 0.013 0.46 $\pm$ 0.006 0.46 $\pm$ 0.006 0.46 $\pm$ 0.006
Arrest Diversion 20, 22, 0 5.18 $\pm$ 0.02 5.17 $\pm$ 0.019 5.19 $\pm$ 0.019 0.66 $\pm$ 0.013∗∗ 0.64 $\pm$ 0.014∗∗ 0.65 $\pm$ 0.013∗∗ 0.51 $\pm$ 0.006∗∗ 0.49 $\pm$ 0.006∗∗ 0.51 $\pm$ 0.006∗∗
40, 22, 0 5.19 $\pm$ 0.021 5.2 $\pm$ 0.021 5.18 $\pm$ 0.02 0.55 $\pm$ 0.012∗∗ 0.54 $\pm$ 0.012∗∗ 0.56 $\pm$ 0.012∗∗ 0.55 $\pm$ 0.006∗∗ 0.56 $\pm$ 0.006∗∗ 0.56 $\pm$ 0.006∗∗
60, 22, 0 5.19 $\pm$ 0.021 5.21 $\pm$ 0.02 5.19 $\pm$ 0.019 0.47 $\pm$ 0.011∗∗ 0.44 $\pm$ 0.011∗∗ 0.44 $\pm$ 0.01∗∗ 0.6 $\pm$ 0.006∗∗ 0.59 $\pm$ 0.006∗∗ 0.6 $\pm$ 0.005∗∗
80, 22, 0 5.22 $\pm$ 0.02 5.15 $\pm$ 0.02 5.18 $\pm$ 0.019 0.34 $\pm$ 0.009∗∗ 0.33 $\pm$ 0.01∗∗ 0.35 $\pm$ 0.009∗∗ 0.65 $\pm$ 0.006∗∗ 0.65 $\pm$ 0.006∗∗ 0.65 $\pm$ 0.006∗∗
100, 22, 0 5.14 $\pm$ 0.02 5.15 $\pm$ 0.02 5.19 $\pm$ 0.019 0.23 $\pm$ 0.008∗∗ 0.24 $\pm$ 0.007∗∗ 0.21 $\pm$ 0.007∗∗ 0.68 $\pm$ 0.006∗∗ 0.7 $\pm$ 0.006∗∗ 0.69 $\pm$ 0.006∗∗
Overdose Diversion 0, 30, 0 4.81 $\pm$ 0.02∗∗ 4.79 $\pm$ 0.02∗∗ 4.84 $\pm$ 0.019∗∗ 0.79 $\pm$ 0.015 0.77 $\pm$ 0.014 0.77 $\pm$ 0.013 0.58 $\pm$ 0.006∗∗ 0.58 $\pm$ 0.006∗∗ 0.58 $\pm$ 0.006∗∗
0, 45, 0 4.03 $\pm$ 0.019∗∗ 4.08 $\pm$ 0.018∗∗ 4.05 $\pm$ 0.017∗∗ 0.78 $\pm$ 0.014 0.73 $\pm$ 0.014 0.76 $\pm$ 0.014 0.81 $\pm$ 0.007∗∗ 0.8 $\pm$ 0.007∗∗ 0.82 $\pm$ 0.007∗∗
0, 60, 0 3.3 $\pm$ 0.017∗∗ 3.32 $\pm$ 0.016∗∗ 3.35 $\pm$ 0.015∗∗ 0.76 $\pm$ 0.014 0.76 $\pm$ 0.014 0.78 $\pm$ 0.014 1.03 $\pm$ 0.007∗∗ 1.04 $\pm$ 0.007∗∗ 1.07 $\pm$ 0.007∗∗
0, 75, 0 2.6 $\pm$ 0.014∗∗ 2.59 $\pm$ 0.014∗∗ 2.6 $\pm$ 0.013∗∗ 0.76 $\pm$ 0.015 0.74 $\pm$ 0.014 0.75 $\pm$ 0.013 1.28 $\pm$ 0.008∗∗ 1.27 $\pm$ 0.008∗∗ 1.28 $\pm$ 0.007∗∗
0, 90, 0 1.9 $\pm$ 0.012∗∗ 1.88 $\pm$ 0.012∗∗ 1.87 $\pm$ 0.011∗∗ 0.75 $\pm$ 0.015 0.77 $\pm$ 0.014 0.76 $\pm$ 0.013 1.51 $\pm$ 0.008∗∗ 1.5 $\pm$ 0.008∗∗ 1.51 $\pm$ 0.008∗∗
Case Management 0, 22, 20 5.19 $\pm$ 0.02 5.21 $\pm$ 0.02 5.2 $\pm$ 0.018 0.66 $\pm$ 0.013∗∗ 0.63 $\pm$ 0.013∗∗ 0.65 $\pm$ 0.012∗∗ 0.53 $\pm$ 0.006∗∗ 0.52 $\pm$ 0.006∗∗ 0.53 $\pm$ 0.005∗∗
0, 22, 40 5.18 $\pm$ 0.021 5.17 $\pm$ 0.019 5.17 $\pm$ 0.02 0.54 $\pm$ 0.012∗∗ 0.53 $\pm$ 0.011∗∗ 0.54 $\pm$ 0.011∗∗ 0.67 $\pm$ 0.006∗∗ 0.66 $\pm$ 0.006∗∗ 0.66 $\pm$ 0.006∗∗
0, 22, 60 5.18 $\pm$ 0.022 5.18 $\pm$ 0.02 5.17 $\pm$ 0.019 0.44 $\pm$ 0.01∗∗ 0.42 $\pm$ 0.01∗∗ 0.43 $\pm$ 0.01∗∗ 0.84 $\pm$ 0.006∗∗ 0.83 $\pm$ 0.006∗∗ 0.83 $\pm$ 0.006∗∗
0, 22, 80 5.17 $\pm$ 0.021 5.13 $\pm$ 0.019 5.16 $\pm$ 0.019 0.3 $\pm$ 0.009∗∗ 0.3 $\pm$ 0.008∗∗ 0.3 $\pm$ 0.008∗∗ 1.03 $\pm$ 0.007∗∗ 1.01 $\pm$ 0.007∗∗ 1.02 $\pm$ 0.007∗∗
0, 22, 100 5.13 $\pm$ 0.02 5.17 $\pm$ 0.021 5.17 $\pm$ 0.02 0.2 $\pm$ 0.007∗∗ 0.19 $\pm$ 0.007∗∗ 0.19 $\pm$ 0.007∗∗ 1.26 $\pm$ 0.007∗∗ 1.23 $\pm$ 0.007∗∗ 1.22 $\pm$ 0.007∗∗
Policy Mix 20, 40, 20 4.27 $\pm$ 0.02∗∗ 4.33 $\pm$ 0.018∗∗ 4.3 $\pm$ 0.018∗∗ 0.58 $\pm$ 0.013∗∗ 0.58 $\pm$ 0.012∗∗ 0.53 $\pm$ 0.011∗∗ 0.87 $\pm$ 0.007∗∗ 0.85 $\pm$ 0.006∗∗ 0.86 $\pm$ 0.007∗∗
40, 60, 40 3.29 $\pm$ 0.016∗∗ 3.28 $\pm$ 0.017∗∗ 3.29 $\pm$ 0.015∗∗ 0.41 $\pm$ 0.011∗∗ 0.42 $\pm$ 0.01∗∗ 0.4 $\pm$ 0.01∗∗ 1.31 $\pm$ 0.008∗∗ 1.3 $\pm$ 0.007∗∗ 1.3 $\pm$ 0.007∗∗
60, 80, 60 2.33 $\pm$ 0.014∗∗ 2.3 $\pm$ 0.013∗∗ 2.32 $\pm$ 0.013∗∗ 0.29 $\pm$ 0.009∗∗ 0.3 $\pm$ 0.009∗∗ 0.3 $\pm$ 0.009∗∗ 1.76 $\pm$ 0.008∗∗ 1.75 $\pm$ 0.008∗∗ 1.75 $\pm$ 0.007∗∗
@@ 3.5 Discussion
3.6 Model Sensitivity
To assess model sensitivity, we used the (60, 80, 60) scenario and sampled 1024 Sobol sets across 41 parameters, ranging each parameter $\pm$ 5% of their base model value. We ran each set with three replications for a total of 3,072 simulation runs. The sensitivity analysis in Appendix D suggests that this paper’s conclusions are not sensitive to model inputs, since the sensitivity analysis shows that the longer the model is run, the less sensitive the model is to the starting population. Additionally, the relationship between the model inputs and outputs is expected due to the model structure. Care should be taken when estimating distributions for arcs (2)-(6) and (G) since these have the largest effect on base model outputs, particularly $\mu$ for these distributions. Overall, the sensitivity analysis suggests that changes to the model parameters would not affect the conclusions of this paper between the base model and treatment policy scenarios regarding opioid-related arrests, hospital encounters, treatments, and active use.
@@ 3.7 Policy Effect Comparison
To compare policies and assess marginal returns, we evaluate the comparative effects between our three policies AD, OD, and CM, and each of the simulation main outputs: overdose-related deaths, number of hospital encounters, number of non-diverted arrests, number of OUD treatment, and number of individuals in the active state at the end of a given year. We use Ordinary Least Squares (OLS) regression, where the policies are variables against each output in years 2023, 2027, and 2032. Table 6 shows the regression coefficients against each policy and the model outputs across 600 replications for each of the scenarios listed in Tables 4 and 5. The coefficients are interpreted as the improvement in the 1-year, 5-year and 10-year cumulative outcome per a 1% increase of the policy implementation level.
As shown in Table 6, all of the policies are consistent in their direction of influence over model outputs. This means the policies are all negative or all positive for a given year and model output. This means that the policies are complementary in nature and that increasing any policy will likely improve model outputs. We also see that the OD policy has the highest magnitude of coefficients, followed by the CM policy. Therefore, if one could only improve a single policy, assuming the same level of implementation increase in all policies, an improvement in OD would likely have the largest impact, followed by the CM policy and then the AD policy. This makes sense, given the model input parameters, since a larger number of individuals experience a hospital encounter than with the CJS It also makes sense that the CM policy has a larger impact than the AD policy since it can engage any individual in the CJS with past opioid misuse rather than just individuals who are in the CJS for an opioid-related crime.
@@
2023 Opioid-related death 96.42 ($<$ 0.001) -0.03 ($<$ 0.001) -0.0(0.29) -0.02 ($<$ 0.001)
2023 Opioid-related arrest 625.2 ($<$ 0.001) -0.16 ($<$ 0.001) -0.04 ($<$ 0.001) -0.08 ($<$ 0.001)
2023 Hospital encounters 2198.77 ($<$ 0.001) -0.55 ($<$ 0.001) -0.09 ($<$ 0.001) -0.48 ($<$ 0.001)
2023 Treatment 1888.38 ($<$ 0.001) 19.2 ($<$ 0.001) 5.18 ($<$ 0.001) 12.98 ($<$ 0.001)
2023 Active use 17859.79 ($<$ 0.001) -4.82 ($<$ 0.001) -0.85 ($<$ 0.001) -3.29 ($<$ 0.001)
2027 Opioid-related death 499.13 ($<$ 0.001) -0.12 ($<$ 0.001) -0.02 ($<$ 0.001) -0.09 ($<$ 0.001)
2027 Opioid-related arrest 3238.95 ($<$ 0.001) -0.77 ($<$ 0.001) -0.14 ($<$ 0.001) -0.58 ($<$ 0.001)
2027 Hospital encounters 11388.14 ($<$ 0.001) -2.56 ($<$ 0.001) -0.45 ($<$ 0.001) -2.3 ($<$ 0.001)
2027 Treatment 9812.7 ($<$ 0.001) 100.18 ($<$ 0.001) 26.93 ($<$ 0.001) 65.73 ($<$ 0.001)
2027 Active use 92765.62 ($<$ 0.001) -22.67 ($<$ 0.001) -3.96 ($<$ 0.001) -16.05 ($<$ 0.001)
2032 Opioid-related death 1042.9 ($<$ 0.001) -0.26 ($<$ 0.001) -0.05 ($<$ 0.001) -0.17 ($<$ 0.001)
2032 Opioid-related arrest 6736.45 ($<$ 0.001) -1.5 ($<$ 0.001) -0.27 ($<$ 0.001) -1.02 ($<$ 0.001)
2032 Hospital encounters 23670.73 ($<$ 0.001) -4.93 ($<$ 0.001) -0.89 ($<$ 0.001) -3.78 ($<$ 0.001)
2032 Treatment 20451.01 ($<$ 0.001) 209.32 ($<$ 0.001) 56.29 ($<$ 0.001) 135.98 ($<$ 0.001)
2032 Active use 193153.61 ($<$ 0.001) -44.0 ($<$ 0.001) -7.79 ($<$ 0.001) -28.5 ($<$ 0.001)